OpenStudy (anonymous):

34 4/8 + 6 8/10 + 2 1/4 =

OpenStudy (anonymous):

34+6+2 = 42 4/8 + 8/10 + 1/4 lcm = 40 20/40 + 32/40 + 10/40 = 62/40 = 1 22/40 = 1 11/20 total = 42 + 1 11/20 = 43 11/20